What is the next term in the geometric sequence -270, 90, -30?
1
Expert's answer
2019-07-22T03:30:43-0400

"b_1=-270, b_2=90, b_3=-30, b_4=?"

Common ratio: "q= \frac {b_2} {b_1} = \frac {90}{-270} = -\frac{1}{3}," then

"b_4=b_3 \cdot q=-30 \cdot (-\frac {1}{3})= 10"


Answer: "b_4=10"